# relsizeFont size

### v2

@@ Abstract
Sample selection improves the efficiency and effectiveness of machine learning models by providing informative and representative samples. Typically, samples can be modeled as a sample graph, where nodes are samples and edges represent their similarities. Most existing methods are based on local information, such as the training difficulty of samples, thereby overlooking global information, such as connectivity patterns. This oversight can result in suboptimal selection because global information is crucial for ensuring that the selected samples well represent the structural properties of the graph. To address this issue, we employ structural entropy to quantify global information and losslessly decompose it from the whole graph to individual nodes using the Shapley value. Based on the decomposition, we present Structural-Entropy-based sample Selection (SES), a method that integrates both global and local information to select informative and representative samples. SES begins by constructing a $k$NN-graph among samples based on their similarities. It then measures sample importance by combining structural entropy (global metric) with training difficulty (local metric). Finally, SES applies importance-biased blue noise sampling to select a set of diverse and representative samples. Comprehensive experiments on three learning scenarios — supervised learning, active learning, and continual learning — clearly demonstrate the effectiveness of our method.
@@ 1 Introduction
Data budgets that limit sample sizes are pervasive in machine learning applications. For example, researchers and practitioners often face limited annotation and computational resources, necessitating the use of fewer but more informative samples to enhance efficiency. Similarly, in continual learning scenarios (Hou et al., 2019), the memory constraint requires fewer but more representative samples from previous tasks to effectively retain knowledge. Consequently, effective sample selection becomes crucial to improving efficiency and effectiveness in machine learning. It aims to select informative and representative samples from large datasets to accelerate training and enhance the training performance. During selection, samples can be modeled as a sample graph, where nodes are samples and edges represent their similarities. Existing sample selection methods primarily focus on local information, such as training difficulty and node degree (Maharana et al., 2024). Although these methods demonstrate promising performance on many datasets, they overlook the global information inherent in the graph structure. This global information, such as connectivity patterns, captures the structural properties of the whole graph (Leskovec & Faloutsos, 2006) and has been shown to be effective in improving the representativeness of selected samples (Zhang et al., 2023; Yuan et al., 2020; Zhao et al., 2021). Therefore, we aim to incorporate global information into the sample selection process to improve the quality of the selected samples.
@@
The key to incorporating global information is to identify which specific metric(s) can accurately capture the global structure of the sample graph. Li & Pan (2016) propose structural entropy to evaluate the amount of information required to describe a given graph structure. The main feature of this metric is that it is robust and sensitive. First, it remains stable against minor changes like the addition or removal of a few edges. This ensures that the structural entropy reliably reflects the global structure of the graph despite potential noise. Second, it is sensitive to topological changes, especially those affecting connectivity patterns. This is essential for capturing the global structure in response to even small topological changes. These two properties make structural entropy an effective metric for quantifying the global structure and therefore valuable in sample selection. However, existing methods only provide a single value for the whole graph. This presents a challenge in decomposing this metric to the level of individual nodes, limiting its utility for fine-grained, node-level selection.
@@
To address this challenge, we use the Shapley value (Shapley, 1951), a method that fairly decomposes a metric among contributors based on their individual contributions. Specifically, it is calculated by evaluating a node’s marginal contribution to structural entropy when adding this node to each subgraph of the sample graph. This decomposition process is highly time-consuming as it requires the enumeration of all possible subgraphs. To accelerate this, we reformulate the Shapley value for structural entropy, enabling linear-time calculation with respect to the number of edges. Based on this reformulation, we develop a node-level structural entropy metric that effectively measures the importance of nodes in preserving the global structure. Building on the decomposition, we present a Structural-Entropy-based sample Selection (SES) method that integrates both global and local metrics to select informative and representative samples. This method begins by constructing a $k$NN-graph among samples to describe their similarity relationships. Then, it measures sample importance by combining node-level structural entropy (global metric) with training difficulty (local metric). Finally, the importance-biased blue noise sampling method is employed to iteratively select a set of diverse and representative samples.
@@
We validate the effectiveness of our method through comprehensive experiments on three important learning scenarios: supervised learning, active learning, and continual learning. The evaluation covers many tasks, including image classification, text classification, object detection, and visual question answering. The results clearly show that our method consistently improves the state-of-the-art methods across all scenarios and tasks. This indicates that our method of integrating global and local information outperforms existing methods in selecting more informative and representative samples.
@@
  • Node-level structural entropy that quantifies the importance of nodes in preserving the global structure.
  • A structural-entropy-based sample selection method that integrates both global and local metrics to select informative and representative samples [^1] The implementation is available at https://anonymous.4open.science/r/SE-based__\__sample__\__selection-575B/..
  • Experiments in supervised learning, active learning, and continual learning that demonstrate the effectiveness of our method.
@@ 2 Related work
Attribute-based methods rely on the attributes of individual samples. A commonly used attribute is the training difficulty, which is typically assessed from two perspectives: confidence and error. Metrics that measure model confidence include the entropy of the prediction vector (Coleman et al., 2020) and the variance of the predicted probabilities across training epochs (Swayamdipta et al., 2020). Metrics that measure model error include EL2N (Paul et al., 2021), which calculates the $L_{2}$ norm of the error vector, and the Forgetting score (Toneva et al., 2019), which tracks the frequency of misclassifications after initial correct classifications. AUM (Pleiss et al., 2020) combines both perspectives by measuring the confidence for correct classifications and the error for misclassifications. Based on these metrics, several sample selection methods have been developed. One simple yet effective method is selecting the most difficult samples, as they have a larger impact on model performance (Paul et al., 2021). However, this method overlooks easy samples, which are crucial for model training when data budgets are limited (Sorscher et al., 2022). To address this issue, CCS (Zheng et al., 2022) divides the dataset into strata based on training difficulty and performs random sampling within each stratum. InfoBatch (Qin et al., 2023) retains some easy samples and enhances their influence by upscaling their gradient. Another line of work uses the gradient as the attribute and aims to match the average gradient of the selected samples with that of all samples (Mirzasoleiman et al., 2019; Killamsetty et al., 2021). However, these gradients depend on the model’s current state during training, limiting the applicability of the selected samples to other models.
@@
While these methods demonstrate promising performance on many datasets, they often overlook the global information, which is crucial for increasing the representativeness of selected samples (Yuan et al., 2020). Overlooking this global information can lead to suboptimal learning performance. To address this gap, we propose a structural-entropy-based sample selection method that integrates both global and local metrics to select informative and representative samples. The experimental results presented in Sec. LABEL:sub@sec:exp show that our method achieves consistent improvement over existing methods.
@@ 3 Background: Structural Entropy of Graph
Structural entropy evaluates how the nodes and edges in a graph are hierarchically organized to form communities at different levels (Li & Pan, 2016). Thus, it is effective in globally quantifying the community structure of the graph regarding its overall connectivity patterns. The calculation of structural entropy is based on an encoding tree that represents the graph’s hierarchical community structure. In this tree, each node represents a community, and each leaf node corresponds to a graph node that forms a single-node community. With this encoding tree, the entropy is aggregated across communities of different levels, which provides insights into the hierarchical community structure of the graph. Fig. LABEL:sub@fig:codingtree shows an undirected, weighted graph $G$ and its encoding tree $\mathcal{T}$. The entropy is calculated for each non-root node $\alpha$, which considers both intra-community and inter-community connections to reflect the connectivity patterns of its community in the graph. Lower entropy indicates denser intra-community connections and sparser inter-community connections. The intra-community connection, $vol(\alpha)$, is quantified by the total weighted degrees of nodes, while the inter-community connection, $g(\alpha)$, is quantified by the total weight of the edges with exactly one endpoint in the community (outer edges). Given an encoding tree $\mathcal{T}$, the structural entropy of $G$ is the aggregation of the entropy values from all its non-root nodes:
@@
In real-world applications, the encoding tree $\mathcal{T}$ may be unknown. To best capture the hierarchical community structure in such cases, the encoding tree $\mathcal{T}$ is constructed by minimizing the structural entropy. Obtaining an exact solution for the minimization is challenging, so a greedy strategy similar to the Huffman tree construction has been developed (Zhu et al., 2023).
@@ 4 Structural-Entropy-Based Sample Selection
Our sample selection method integrates global and local metrics to select informative and representative samples. Given a large set of samples, an undirected, weighted sample graph $G$ is initially constructed to model their similarity relationships. Each sample, represented by its embedding extracted by a deep neural network, corresponds to a node in the graph. To avoid excessive edge connections, each sample is connected to its $k$ nearest neighboring samples. The edge weight between any two samples, $u$ and $v$, is their cosine similarity normalized to $[0,1]$. Based on this graph, we first propose a node-level structural entropy metric to globally measure the importance of each sample. Then, it is combined with a local metric, training difficulty, to assign an importance score to each sample. Using this score, we develop an importance-biased blue noise sampling method to select a set of informative and representative samples.
@@ 4.1 Node-Level Structural Entropy
Chen & Teng (2017) have shown that Shapley value (Shapley, 1951) is an effective method to decompose a value from the graph level to the node level. The key feature of this method is its lossless and fair decomposition of the value, ensuring that the aggregate node-level value equals the graph-level value. Inspired by this, we employ the Shapley value to derive the node-level structural entropy. Specifically, the Shapley value of a node $u$ reflects the average increase in structural entropy when it is added to all possible subgraphs of $G$. As a result, this value captures the node’s contribution to the global connectivity patterns.
@@
Proof.
The proof of this proposition is provided in Appendix LABEL:sub@appendix:proof. ∎
Proposition LABEL:sub@prop:se shows that the Shapley value can be calculated in linear time with respect to the number of edges. Thus, this reformulation enables an efficient and exact calculation. In particular, the calculation consists of two terms. The first decomposes the structural entropy from the encoding tree to the node, and the second reflects local connectivity through node degree. Given our emphasis on developing a global metric, we only use the first term to define the node-level structural entropy ($S_{e}$):
@@
A higher $vol(\alpha_{u\vee v})$ indicates that the edge $\langle u,v\rangle$ connects nodes that are more distantly located in $\mathcal{T}$. For example, in Fig. LABEL:sub@fig:lca(a), the edge $\langle u,v\rangle$ connects nodes across different communities. In contrast, in Fig. LABEL:sub@fig:lca(b), the edge $\langle u,v^{\prime}\rangle$ connects nodes within the same community, resulting in a lower $vol(\alpha_{u\vee v^{\prime}})$ compared to $vol(\alpha_{u\vee v})$. Thus, $vol(\alpha_{u\vee v})$ effectively quantifies the extent to which an edge bridges different communities. Given that the node-level structural entropy is the weighted sum of $vol(\alpha_{u\vee v})$, nodes with high structural entropy serve as boundaries between communities. Selecting these nodes is crucial for maintaining the overall structure of the graph.
@@ 4.2 Importance-Biased Blue Noise Sampling
To select a set of high-quality samples, the developed global metric, node-level structural entropy, needs to be combined with an appropriate local metric. Previous research has shown that training difficulty ($S_{t}$) is an effective local metric in quantifying the sample’s impact on model performance (Paul et al., 2021; Sorscher et al., 2022). Based on this, we employ it as the local metric. Accordingly, the overall importance score ($S$) of a sample is a combination of node-level structural entropy and training difficulty:
@@
Given the importance scores, a straightforward solution is to select the samples with the highest scores. However, this significantly reduces the diversity of the selected samples, as the important samples tend to cluster in several narrow regions (Zheng et al., 2022). An alternative is the message passing mechanism employed by $\mathbb{D}^{2}$ Pruning: once a sample is selected, this method sends weighted messages to decrease the importance scores of its neighbors in the graph. However, the message weights are sensitive to a hyperparameter and can lead to suboptimal results if not carefully tuned. Previous research has shown that blue noise sampling achieves a good balance between randomness and uniformity by excluding overly similar samples (Xiang et al., 2019; Liu et al., 2017). This method increases sampling in low-density regions, which enhances the diversity of the selected samples. Consequently, we develop an importance-biased blue noise sampling method to select a set of informative and representative samples.
Our sampling process contains two steps: 1) identifying the candidate sample with the highest importance score, 2) rejecting the sample if its similarity with any selected neighboring samples exceeds a threshold $\theta$; otherwise, accepting it as a selected sample. These two steps are iteratively performed until no more samples can be selected. To determine the threshold $\theta$ for a given sampling rate, we perform a binary search on $\theta$.

### v3

@@ Abstract
Sample selection improves the efficiency and effectiveness of machine learning models by providing informative and representative samples. Typically, samples can be modeled as a sample graph, where nodes are samples and edges represent their similarities. Most existing methods are based on local information, such as the training difficulty of samples, thereby overlooking global information, such as connectivity patterns. This oversight can result in suboptimal selection because global information is crucial for ensuring that the selected samples well represent the structural properties of the graph. To address this issue, we employ structural entropy to quantify global information and losslessly decompose it from the whole graph to individual nodes using the Shapley value. Based on the decomposition, we present Structural-Entropy-based sample Selection (SES), a method that integrates both global and local information to select informative and representative samples. SES begins by constructing a $k$NN-graph among samples based on their similarities. It then measures sample importance by combining structural entropy (global metric) with training difficulty (local metric). Finally, SES applies importance-biased blue noise sampling to select a set of diverse and representative samples. Comprehensive experiments in three learning scenarios — supervised learning, active learning, and continual learning — clearly demonstrate the effectiveness of our method.
@@ 1 Introduction
Data budgets that limit sample sizes are pervasive in machine learning applications (Liu et al., 2025). For example, researchers and practitioners often face limited annotation and computational resources, necessitating the use of fewer samples to enhance efficiency. Similarly, in continual learning scenarios (Hou et al., 2019), the memory constraint requires fewer samples from previous tasks to retain knowledge effectively. Consequently, effective sample selection becomes crucial to improving efficiency and effectiveness in machine learning. It aims to select informative and representative samples from large datasets to accelerate training and enhance the training performance. Informative samples are those that significantly reduce model uncertainty and are crucial for improving the accuracy and robustness of the training process, while representative samples are those that preserve the diversity and overall distribution of the dataset (Huang et al., 2014). During selection, samples can be modeled as a sample graph, where nodes are samples and edges represent their similarities. Existing sample selection methods primarily focus on local information, such as the training difficulty and the node degree (Maharana et al., 2024). Although these methods demonstrate promising performance on many datasets, they overlook the global information inherent in the graph structure. This global information, such as connectivity patterns, captures the structural properties of the whole graph (Leskovec & Faloutsos, 2006) and has been shown to be effective in improving the representativeness of selected samples (Zhang et al., 2023; Yuan et al., 2021; Zhao et al., 2021). Therefore, we aim to incorporate global information into the sample selection process to improve the quality of the selected samples.
@@
The key to incorporating global information is to identify which specific metric(s) can accurately capture the global structure of the sample graph. Entropy is a class of metrics well-suited for this purpose, as it quantifies both informativeness and representativeness (Pan et al., 2005; Li & Guo, 2013). In particular, Li & Pan (2016) propose structural entropy to evaluate the amount of information required to describe a given graph structure. The main feature of this metric is that it is robust and sensitive. First, it remains stable against minor changes like the addition or removal of a few edges. This ensures that the structural entropy reliably reflects the global structure of the graph despite potential noise. Second, it is sensitive to topological changes, especially those affecting connectivity patterns. This is essential for capturing the global structure in response to even small topological changes. These two properties make structural entropy an effective metric for quantifying the global structure and therefore valuable in sample selection. However, existing methods only provide a single value for the whole graph. This presents a challenge in decomposing this metric to the level of individual nodes, limiting its utility for fine-grained, node-level selection.
@@
To address this challenge, we use the Shapley value (Shapley, 1951), a method that fairly decomposes a metric among contributors based on their individual contributions. Specifically, it is calculated by evaluating a node’s marginal contribution to structural entropy when adding this node to each subgraph of the sample graph. This decomposition process is highly time-consuming as it requires exponential-time computation to enumerate possible subgraphs. To accelerate this, we reformulate the Shapley value for structural entropy, enabling linear-time calculation with respect to the edge number. Based on this reformulation, we propose a node-level structural entropy metric that effectively measures the importance of nodes in preserving the global structure. Building on the decomposition, we present a Structural-Entropy-based sample Selection (SES) method that integrates both global and local metrics to select informative and representative samples. This method begins by constructing a $k$NN-graph among samples to describe their similarity relationships. Then, it measures sample importance by combining node-level structural entropy (global metric) with training difficulty (local metric). Finally, the importance-biased blue noise sampling method is employed to iteratively select a set of diverse and representative samples.
@@
We validate the effectiveness of our method through comprehensive experiments on three important learning scenarios: supervised learning, active learning, and continual learning. The evaluation covers many tasks, including image classification, text classification, object detection, and visual question answering. The results clearly show that our method consistently improves state-of-the-art methods across all scenarios and tasks. This indicates that our method of integrating global and local information outperforms existing methods in selecting more informative and representative samples.
@@
  • We propose a node-level structural entropy metric that quantifies the importance of nodes in preserving the global structure, and it can be calculated in linear time.
  • We develop a structural-entropy-based sample selection method that integrates both global and local metrics to select informative and representative samples [^1] The implementation is available at https://github.com/thu-vis/SE-based__\__sample__\__selection..
  • We conduct experiments in supervised learning, active learning, and continual learning that demonstrate the effectiveness of our method.
@@ 2 Related work
Attribute-based methods rely on the attributes of individual samples. A commonly used attribute is the training difficulty, which is typically assessed from two perspectives: confidence and error. Metrics that measure model confidence include the entropy of the prediction vector (Coleman et al., 2020) and the variance of the predicted probabilities across training epochs (Swayamdipta et al., 2020). Metrics that measure model error include EL2N (Paul et al., 2021), which calculates the $L_{2}$ norm of the error vector, and the Forgetting score (Toneva et al., 2019), which tracks the frequency of misclassifications after initial correct classifications. AUM (Pleiss et al., 2020) combines both perspectives by measuring the confidence for correct classifications and the error for misclassifications. Based on these metrics, several sample selection methods have been developed. One simple yet effective method is selecting the most difficult samples, as they have a larger impact on the model performance (Paul et al., 2021). However, this method overlooks easy samples, which are crucial for model training when data budgets are limited (Sorscher et al., 2022). To address this issue, CCS (Zheng et al., 2022) divides the dataset into strata based on training difficulty and performs random sampling within each stratum. InfoBatch (Qin et al., 2023) retains some easy samples and enhances their influence by upscaling their gradient. Another line of work uses the gradient as the attribute and aims to match the average gradient of the selected samples with that of all samples (Mirzasoleiman et al., 2019; Killamsetty et al., 2021). However, these gradients depend on the model’s current state during training, limiting the applicability of the selected samples to other models.
@@
While these methods demonstrate promising performance on many datasets, they often overlook the global information, which is crucial for increasing the representativeness of selected samples (Yuan et al., 2021). Overlooking this global information can lead to suboptimal learning performance. To address this gap, we propose a structural-entropy-based sample selection method that integrates both global and local metrics to select informative and representative samples.
@@ 3 Background: Structural Entropy of Graph
Structural entropy evaluates how the nodes and edges in a graph are hierarchically organized to form communities at different levels (Li & Pan, 2016). Thus, it is effective in globally quantifying the community structure of the graph regarding its overall connectivity patterns. The calculation of structural entropy is based on an encoding tree that represents the graph’s hierarchical community structure. In this tree, each node represents a community, and each leaf node corresponds to a graph node that forms a single-node community. With this encoding tree, the entropy is aggregated across communities of different levels, which provides insights into the hierarchical community structure of the graph. Fig. LABEL:sub@fig:codingtree shows an undirected, weighted graph $G$ and its encoding tree $\mathcal{T}$. The entropy is calculated for each non-root node $\alpha$, which considers both intra-community and inter-community connections to reflect the connectivity patterns of its community in the graph. Lower entropy indicates denser intra-community connections and sparser inter-community connections. The intra-community connection, $vol(\alpha)$, is quantified by the total weighted degrees of the nodes, while the inter-community connection, $g(\alpha)$, is quantified by the total weight of the edges with exactly one endpoint in the community (outer edges). Given an encoding tree $\mathcal{T}$, the structural entropy of $G$ is the aggregation of the entropy values from all its non-root nodes:
@@
In real-world applications, the encoding tree $\mathcal{T}$ may be unknown. To best capture the hierarchical community structure in such cases, the encoding tree $\mathcal{T}$ is constructed by minimizing the structural entropy. Obtaining an exact solution for the minimization is challenging, so greedy methods similar to the Huffman tree construction have been developed (Li & Pan, 2016; Zhu et al., 2023). We demonstrate in Appendix LABEL:sub@appendix:SE_construction that the choice of the construction method has a negligible effect on the sample selection results. Therefore, we use the most recent method proposed by Zhu et al. (2023).
@@ 4 Structural-Entropy-Based Sample Selection
Our sample selection method integrates global and local metrics to select informative and representative samples. Given a large set of samples, an undirected, weighted sample graph $G$ is initially constructed to model their similarity relationships. Each sample, represented by its embedding extracted by a deep neural network, corresponds uniquely to a node in the graph. To avoid excessive edge connections, each sample is connected to its $k$ nearest neighboring samples. The edge weight between any two samples, $u$ and $v$, is their cosine similarity normalized to $[0,1]$. Based on this graph, we first propose a node-level structural entropy metric to globally measure the importance of each sample. Then, it is combined with a local metric, training difficulty, to assign an importance score to each sample. Using this score, we develop an importance-biased blue noise sampling method to select a set of informative and representative samples.
@@ 4.1 Node-Level Structural Entropy
Chen & Teng (2017) have shown that the Shapley value (Shapley, 1951) is an effective method to decompose a value from the graph level to the node level. The key feature of this method is its lossless and fair decomposition of the value, ensuring that the aggregate node-level value equals the graph-level value. Inspired by this, we employ the Shapley value to derive the node-level structural entropy. Specifically, the Shapley value of a node $u$ reflects the average increase in structural entropy when it is added to all possible subgraphs of $G$. As a result, this value captures the node’s contribution to the global connectivity patterns.
@@
The proof of Proposition LABEL:sub@prop:se is provided in Appendix LABEL:sub@appendix:proof. It indicates that the Shapley value can be calculated in linear time with respect to the edge number. Thus, this reformulation enables an efficient and exact calculation. Eq. (LABEL:sub@equ:shapley_result) consists of two terms. The first decomposes the structural entropy from the encoding tree to the node, and the second reflects local connectivity through node degree. Due to the below theoretical and empirical advantages, we only use the first term to define the node-level structural entropy ($S_{e}$):
@@
[…]
$\frac{1}{nk^{2}R}\exp\left(\frac{1}{kR}\mathbb{E}[S_{e}(u)]\right)\leq\mathbb{E}[P(u,r)],\$ (5)
where $n$ is the number of samples, $k$ is the parameter in $k$NN-graph construction, and $R$ is the upper bound of edge weights. This result indicates that maximizing $S_{e}(u)$ during selection inherently improves sample coverage. Given the strong correlation between coverage ability and the empirical loss (Zheng et al., 2022), selecting samples with high node-level structural entropy enhances model performance. The mathematical proof and the empirical validation are detailed in Appendix LABEL:sub@appendix:proof2.
The second theoretical advantage is its ability in maintaining the overall graph structure. A higher $vol(\alpha_{u\vee v})$ indicates that the edge $\langle u,v\rangle$ connects the nodes that are more distantly located in $\mathcal{T}$. For example, in Fig. LABEL:sub@fig:lca, the edge $\langle u,v^{\prime}\rangle$ stays within the same community while the edge $\langle u,v\rangle$ spans different communities, resulting in a higher $vol(\alpha_{u\vee v})$. Thus, $vol(\alpha_{u\vee v})$ effectively quantifies the extent to which an edge bridges different communities. Since $S_{e}(u)$ is the weighted sum of $vol(\alpha_{u\vee v})$, nodes with high structural entropy serve as boundaries between communities. Selecting these nodes is crucial for maintaining the overall structure of the graph.
The empirical advantage of this definition is demonstrated through an ablation study, as detailed in Appendix LABEL:sub@appendix:add_ablat. The results show that only using the first term slightly improves performance.
@@ 4.2 Importance-Biased Blue Noise Sampling
To select a set of high-quality samples, the developed global metric, node-level structural entropy, needs to be combined with an appropriate local metric. Previous research has shown that training difficulty ($S_{t}$) is an effective local metric in quantifying the sample’s impact on model performance, as difficult samples are typically more informative for improving the decision boundary (Paul et al., 2021; Sorscher et al., 2022). Therefore, we employ it as the local metric. Accordingly, the overall importance score ($S$) is a combination of node-level structural entropy and training difficulty:
@@
Given the importance scores, a straightforward solution is to select the samples with the highest scores. However, this significantly reduces the diversity of the selected samples, as the important samples tend to cluster in several narrow regions (Zheng et al., 2022). An alternative is the message passing mechanism employed by $\mathbb{D}^{2}$ Pruning: once a sample is selected, this method sends weighted messages to decrease the importance scores of its neighbors in the graph. However, the message weights are sensitive to a hyperparameter and can lead to suboptimal results if not carefully tuned. Previous research has shown that blue noise sampling achieves a good balance between randomness and uniformity by excluding overly similar samples (Xiang et al., 2019; Liu et al., 2018). This method increases sampling in low-density regions, which enhances the diversity of the selected samples. Consequently, we develop an importance-biased blue noise sampling method to select a set of informative and representative samples.
Our sampling process contains two steps: 1) identifying the candidate sample with the highest importance score, 2) rejecting the sample if its similarity with any selected neighboring samples exceeds a threshold $\theta$; otherwise, accepting it as a selected sample. These two steps are performed iteratively until no more samples can be selected. To determine the threshold $\theta$ for a given sampling rate, we perform a binary search on $\theta$.